OpenStudy (anonymous):

NEED HELP. show that each of the following statements is an identity by transforming the left side of each one into the right side. cos θ tan θ = sin θ

OpenStudy (anonymous):

Rewrite \(\tan\theta\) as \(\dfrac{\sin\theta}{\cos\theta}\) and then cancel out the cosines.